Lakeside Inc. manufactures four lines of remote control boats and uses activity-based costing to calculate product cost.

Activity Cost Pools Estimated Total Cost Estimated Cost Driver
Machining $ 318,000 13,800 machine hours
Setup 65,000 450 batches
Quality control 96,000 790 inspections

Suppose the Luxury boat requires 3,850 machine hours, 115 inspections, and 65 batches. Using activity proportions, determine the amount of overhead assigned to the Luxury product line. (Round your intermediate calculations to the nearest whole percentage and your final answer to the nearest whole number.)

Allocation overhead for maching, setup cost, quality control, total overhead
Answer #1

SOLUTION

Cost per machine hour = 318,000 / 13,800 = 23.04 per machine hour

Cost per batch = 65,000 / 450 = 144.44 per batch

Cost per inspection = 96,000 / 790 = 121.52 per inspection

Cost pool Allocated overheads
Machining (23.04*3,850) = 88,704
Setup cost (144.44*65) = 9,389
Quality control (121.52*115) = 13,975
Total overheads 112,068